# TabNSA: Native Sparse Attention for Efficient Tabular Data Learning thanks: Code on GitHub: https://github.com/aseslamian/TabNSA

### v1

@@ IV-F Regression Task
To evaluate model performance on a regression task, we compared five methods—MLP, XGBoost, CatBoost, TabNet, and our proposed TabNSA—using Root Mean Squared Error (RMSE) as the metric across three datasets: SA, TO, and CP. Lower RMSE indicates better performance. As shown in Table IV, XGBoost achieved the best overall performance, with the lowest RMSE on the CP dataset and competitive scores across the others, resulting in the highest ranking. CatBoost closely followed, outperforming XGBoost on the SA dataset and ranking second overall. The MLP model struggled with relatively higher RMSE values, while TabNet exhibited significantly worse performance, particularly on the SA dataset. Our TabNSA model achieved the lowest RMSE on the TO dataset, demonstrating its effectiveness in certain cases. However, it reported a higher RMSE on the CP dataset, impacting its overall ranking. These results highlight the effectiveness of gradient-boosting models for tabular regression, while also showcasing TabNSA’s strengths in specific domains.
@@
TABLE IV: Evaluation of Different Models for Regression task. Results shows RMSE
@@
MLP 0.607 0.037 7.895 4
XGBoost 0.185 0.0276 2.631 1.67
CatBoost 0.183 0.0281 2.848 2
TabNet 3.2450 0.1050 3.405 4.33
TabNSA 0.3129 0.0273 14.2931 3
V Conclusion
Our experiments demonstrate TabNSA’s superior performance on classification tasks, while showing promising results on certain regression datasets that suggest potential for further improvement. TabNSA’s advantages over previous Transformer-based approaches include reduced computational complexity, hardware-aligned design, and native integration of sparsity during training. These properties enable the model to effectively balance global context with local feature interactions.
TabNSA features reduced computational complexity, hardware-aligned design, and inherent sparsity during training, allowing it to effectively balance global context and local feature interactions.

### v2

@@ Appendix B Regression Task
To evaluate model performance on a regression task, we compared six methods—MLP, XGBoost, CatBoost, TabNet, NSA only and our proposed TabNSA—using Root Mean Squared Error (RMSE) as the metric across three datasets: SA, TO, and CP. Lower RMSE indicates better performance. As shown in Table 13, XGBoost achieved the best overall performance, with the lowest RMSE on the CP dataset and competitive scores across the others, resulting in the highest ranking. CatBoost closely followed, outperforming XGBoost on the SA dataset and ranking second overall. The MLP model struggled with relatively higher RMSE values, while TabNet exhibited significantly worse performance, particularly on the SA dataset. Our TabNSA model achieved the lowest RMSE on the TO dataset, demonstrating its effectiveness in certain cases. However, it reported a higher RMSE on the CP dataset, impacting its overall ranking. These results highlight the effectiveness of gradient-boosting models for tabular regression, while also showcasing TabNSA’s strengths in specific domains.
@@
Table 13: Evaluation of Different Models for Regression task. Results shows RMSE
@@
MLP 0.607 0.037 7.895 5
XGBoost 0.185 0.0276 2.631 2
CatBoost 0.183 0.0281 2.848 2.33
TabNet 3.2450 0.1050 3.405 5.33
NSA 0.3129 0.0273 14.2931 3.67
TabNSA 0.1781 0.0291 2.9814 2.67
  • Note: While TabNSA performs competitively, tree-based models remain superior for regression, highlighting opportunities for further model refinement.
As shown in Table 13, TabNSA performs competitively but does not consistently surpass ensemble-based models like XGBoost and CatBoost on regression benchmarks. This gap may arise from architectural factors: TabNSA’s discrete attention and feature selection mechanisms are well-suited to classification tasks, where identifying sparse, instance-specific relevant features is critical. However, regression problems often involve smoother, continuous relationships between features and outputs, which may not benefit as directly from hard or coarse feature selection. Furthermore, the inductive bias of MLP-style backbones and sparse attention may underfit certain continuous-response distributions compared to tree-based ensembles that naturally partition the input space with high flexibility. Addressing this limitation is an important direction for future work. Possible improvements include incorporating soft attention over continuous-valued targets, hybridizing TabNSA with kernel-based output layers, or using continuous feature embeddings with uncertainty-aware attention mechanisms to better model real-valued outputs. Future work will explore adapting TabNSA for regression by incorporating soft-attention or hybrid continuous embedding mechanisms.
Appendix C Dataset
We provide links to publicly available datasets used for benchmarking in this study. These datasets enable a comprehensive assessment of TabNSA’s performance across various sectors and data types. Summary details are presented in Table 2, with further descriptions provided below.
Financial Decision-Making:
Credit-g (CG) and Credit-Approval (CA) datasets, focused on credit risk evaluation and loan approval scenarios.
Retail:
Dresses-sale (DS) dataset, which includes detailed transaction records from dress sales.
Demographic Analysis:
Adult (AD) and 1995-Income (IO) datasets, covering income classification and census data.
Specialized Industries:
Cylinder Bands (CB) dataset, representing data from the manufacturing sector. Blastchar (BL) dataset, related to materials science. Insurance Company (IC) dataset, offering insights into the insurance domain.
Blood dataset contains data from a blood transfusion service in Taiwan, with 4 attributes describing 748 donors. The task is to predict whether a donor returns for another donation, with 178 positive cases.
Bank dataset includes 45,211 records from a direct marketing campaign by a Portuguese bank, with 16 features per record. The target is to predict term deposit subscriptions, with 5,289 positive labels.
Credit-g dataset describes 1,000 German credit applicants using 20 attributes. The binary classification task is to assess creditworthiness, identifying 700 individuals as having good credit.
Heart dataset aggregates data from four hospitals, comprising 918 patient records with 11 clinical features each. The goal is to detect coronary artery disease, with 508 positive cases.
Income (Adult) dataset consists of 48,842 records from the 1994 U.S. Census, each with 12 attributes. The task is to predict whether an individual’s annual income exceeds \$50,000, with 11,687 positive examples.
Jungle dataset contains 44,819 end-game positions from Jungle Chess, used for specialized evaluation scenarios.
Table 14: Benchmark Dataset Links
Dataset URL
Credit-G https://www.openml.org/search?type=data&status=active&id=31
Credit-Approval https://archive.ics.uci.edu/ml/datasets/credit+approval
Dress-Sales https://www.openml.org/search?type=data&status=active&id=23381
Adult https://www.openml.org/search?type=data&status=active&id=1590
Cylinder-Bands https://www.openml.org/search?type=data&status=active&id=6332
Blastchar https://www.kaggle.com/datasets/blastchar/telco-customer-churn
Insurance-Co https://archive.ics.uci.edu/ml/datasets/Insurance+Company+Benchmark+(COIL+2000)
1995-Income https://www.kaggle.com/datasets/lodetomasi1995/income-classification
Bank https://www.openml.org/search?type=data&status=active&id=45065
Blood https://www.openml.org/search?type=data&status=active&id=1464
Diabetes https://www.kaggle.com/datasets/uciml/pima-indians-diabetes-database
Heart https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction
Jungle https://www.openml.org/search?type=data&status=active&id=41027
ImageSegmentation https://www.openml.org/search?type=data&status=active&id=36
ForestCovertype https://www.openml.org/search?type=data&status=active&id=150
Appendix D Sensivity Analysis
The provided plot of AUC versus the number of Optuna trials illustrates the sensitivity of TabNSA’s performance to the hyperparameter search budget. The curve shows a sharp increase in performance within the initial 20-40 trials, after which the rate of improvement diminishes significantly. By approximately 60-100 trials, the performance has fully plateaued, indicating that further searching yields negligible gains.
This trajectory demonstrates two key findings. First, Optuna’s Bayesian sampler efficiently converges towards optimal hyperparameter configurations, rapidly capturing the model’s inductive biases. Second, and more practically, this analysis reveals that a modest budget of approximately 50 trials is sufficient to achieve over 95% of the model’s peak performance. This provides a clear and practical guideline for researchers and practitioners seeking to apply TabNSA effectively without incurring prohibitive computational costs.
Figure 7: Search Budget Sensitivity Curve: The figure shows the best validation AUC achieved by TabNSA as a function of the number of Optuna trials, across three datasets (DS, CG, and CA). The curve illustrates how model performance improves with increased hyperparameter search budget.
Appendix E Component-wise ablation of sparse attention parameters
We conducted a component-wise ablation study to analyze each sparse attention parameter’s contribution to model performance. The results can be seen in Fig. 8. Specifically, we varied one parameter at a time while keeping others fixed and evaluated the results on four validation sets. For compress block size, a value of 10 achieved the highest average validation AUC ($\sim$0.87), with performance declining at smaller or larger values. Selecting 3 blocks for the number of selected blocks yielded optimal performance ($\sim$0.86), while fewer or more blocks reduced effectiveness. A selection block size of 10 was optimal ($\sim$0.86), with deviations in either direction slightly lowering performance. Lastly, a sliding window size of 6 delivered the best performance ($\sim$0.88), demonstrating the importance of capturing appropriate local context. With this information, we can constrain the search space in the first-stage optimization to accelerate convergence toward the optimal AUC. However, these hyperparameters should still be searched jointly to effectively optimize the model for each dataset.
(a) Impact of selected blocks on Validation Performance
(b) Impact of selection block size on Validation Performance
(c) Impact of compress block size on Validation Performance
(d) Impact of sliding window size on Validation Performance
Figure 8: Component-wise ablation of sparse attention parameters. Each plot shows the effect of varying one parameter : number_of_selected blocks (a), selection_block_size (b), compress_block_size (c) and sliding_window_size (d) on the average validation AUC across four datasets. The results highlight the importance of careful tuning and indicate the optimal value ranges for each parameter.